# Kernel Three Pass Regression Filter

### v1

@@ Abstract
We forecast a single time series using a high-dimensional set of predictors. When these predictors share common underlying dynamics, an approximate latent factor model provides a powerful characterization of their co-movements Bai, (2003). These latent factors succinctly summarize the data and can also be used for prediction, alleviating the curse of dimensionality in high-dimensional prediction exercises, see Stock & Watson, (2002a). However, forecasting using these latent factors suffers from two potential drawbacks. First, not all pervasive factors among the set of predictors may be relevant, and using all of them can lead to inefficient forecasts. The second shortcoming is the assumption of linear dependence of predictors on the underlying factors. The first issue can be addressed by using some form of supervision, which leads to the omission of irrelevant information. One example is the three-pass regression filter proposed by Kelly & Pruitt, (2015). We extend their framework to cases where the form of dependence might be nonlinear by developing a new estimator, which we refer to as the Kernel Three-Pass Regression Filter (K3PRF). This alleviates the aforementioned second shortcoming. The estimator is computationally efficient and performs well empirically. The short-term performance matches or exceeds that of established models, while the long-term performance shows significant improvement.
Keywords: Forecasting, High dimension, Approximate factor model, Reproducing Kernel Hilbert space, Three-pass regression filter.
@@ 1 Introduction
In recent years, high-dimensional datasets have become increasingly accessible across various fields, including economics. It is widely acknowledged that estimation in high-dimensional settings poses a significant challenge known as the ‘curse of dimensionality’, which renders traditional finite-dimensional approaches ineffective.
@@
Most modeling techniques applied to high-dimensional data assume the existence of a low-dimensional structure that effectively summarizes the data. One stylized feature of high-dimensional economic datasets is the presence of high and pervasive collinearity among variables, leading researchers to posit a data-generating process that assumes all variables are a function of a few latent factors. This formulation is commonly referred to as the factor model. There is a vast literature which focuses on using this latent factor structure for forecasting applications. A typical example is found in diffusion index models (Stock & Watson, (2002b)), where latent factors are derived from a high-dimensional set of variables using Principal Components Analysis (hereafter, PCA). These factors are subsequently utilized to forecast a target variable. A limitation of this PCA-based factor estimation is its unsupervised nature, i.e. that no information from the target variable is incorporated during factor estimation.
@@
Given that the primary goal is to forecast a target rather than estimate the underlying factor structure, introducing a degree of supervision can prove beneficial. This can help in filtering out irrelevant information from the predictor set, thus enhancing the predictive accuracy .This can be done in different ways; using soft and hard thresholding methods to remove predictors with no predictive content, as in Bai & Ng, (2008), or assign varying weights to predictors based on their predictive capabilities for the target (see, for example, Huang et al., (2022)), or estimate the subset of factors that exhibit predictive power for the target rather than the complete set of factors that drive the target, as in Kelly & Pruitt, (2015).
@@
The aforementioned models, whether utilizing PCA or supervised factor models, are predicated on the convenient assumption of linearity. However, as underscored in Goulet Coulombe et al., (2022), non-linearity often characterizes many predictive relationships, particularly over extended time horizons and within data-rich environments.
@@
Various approaches have been proposed to integrate non-linearity into factor models. For instance, squared principal components (PCs) or principal component squared ($PC^{2}$) as seen in Bai & Ng, (2008), sufficient forecasting by Fan et al., (2017), the kernel trick to estimate factors (Kutateladze, (2022)) among others. However, these approaches have limited supervision in the prediction process, if any. For example, Fan et al., (2017) estimates factors through an unsupervised method (PCA) and then derives sufficient indices using these PCs. Similarly, Kutateladze, (2022) essentially applies kernel PCA to estimate the set of factors driving a higher-dimensional space obtained by lifting the set of predictors through the kernel method. Similarly, Bai & Ng, (2008), despite employing thresholding to reduce squared PCs to a smaller set, may still encounter challenges in potentially estimating irrelevant factors within this subset, leading to inefficiency. Moreover, in Bai & Ng, (2008), a very particular form of non-linearity (quadratic) is examined, which is somewhat ad hoc.
@@
Our paper incorporates non-linearity by introducing a novel kernel three-pass regression filter. Our approach essentially applies the three-pass filter (hereafter 3PRF) proposed by Kelly & Pruitt, (2015) to a transformed set of predictors. We adopt the lifting concept similar to Kutateladze, (2022), but instead of employing an unsupervised method like kernel PCA, we utilize a supervised method to estimate factors relevant to the target variable. The basic idea behind our approach is that many non-linear relationships observed in one space exhibit linearity in another space.
@@
The table below summarizes our discussion by listing some popular methods in literature and how this paper is placed among them.
Table 1: Factor Model Based Forecasting Methods
@@
The paper proceeds as follows. Section 2 provides a brief introduction to Kernel methods and lists examples of some widely used kernels. Section 3 introduces our estimator and discusses its similarity with the estimator of Kelly & Pruitt, (2015). We also list a set of assumptions that ensure the theoretical properties our estimators, which are given in the subsequent section 4. We present our empirical results in sections 5 and 6 and we finally conclude in section 6. Mathematical proofs and implementation details are given in the appendix.
@@
We use $\boldsymbol{y}$ to denote the $\text{T}\times 1$ vector of the target variable, i.e. $\boldsymbol{y}=(y_{h},y_{h+1}\dots y_{t+h})$. We have $N$ predictors with $T$ observations for each predictor. The cross section of predictors at a time $t$ is given by the $N\times 1$ vector $\boldsymbol{x}_{t}$. Similarly, the vector of temporal observations of a predictor $i$ is given by $\mathbf{x}_{i}$. We stack the predictors in a ${T}\times{N}$ matrix $\boldsymbol{X}$, $\boldsymbol{X}=$ $\left(\boldsymbol{x}_{1}^{\prime},\boldsymbol{x}_{2}^{\prime},\ldots,\boldsymbol{x}_{T}^{\prime}\right)^{\prime}=\left(\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{N}\right)$. We have $L$ proxies which we stack in a ${T}\times\ {L}$ matrix $\boldsymbol{Z}=\left(\boldsymbol{z}_{1}^{\prime},\boldsymbol{z}_{2}^{\prime},\ldots,\boldsymbol{z}_{T}^{\prime}\right)^{\prime}$ $\boldsymbol{J}_{T}\equiv\boldsymbol{I}_{T}-\frac{1}{T}\mathbf{\iota}_{T}\mathbf{\iota}_{T}^{\prime}$, where $\boldsymbol{I}_{T}$ is the $T$-dimensional identity matrix and $\mathbf{\iota}_{T}$ the $T$-vector of ones. For matrices $\boldsymbol{U}$ and $\boldsymbol{V}$ of conformable dimensions, $\boldsymbol{S}_{UV}\equiv\boldsymbol{U}^{\prime}\boldsymbol{J}_{T}\boldsymbol{V}$. For the transformed set of predictors $\varphi(\boldsymbol{X})$, $\varphi_{j}(\mathbf{x})$ denotes the $j^{th}$ observation. $\varphi(\boldsymbol{X})=$ $\left(\varphi_{1}(\mathbf{x}),\varphi_{2}(\mathbf{x}),\ldots,\varphi_{M}(\mathbf{x})\right)=$ $\left(\varphi(\boldsymbol{x}_{1})^{\prime},\varphi(\boldsymbol{x}_{2})^{\prime},\ldots,\varphi(\boldsymbol{x}_{T})^{\prime}\right)^{\prime}$ Stochastic orders are denoted by the usual ${O}_{p}$ and ${o}_{p}$. For a matrix, $\boldsymbol{O}_{p}$ and $\boldsymbol{o}_{p}$ denotes the element wise stochastic order, i.e., a matrix is said to be $\boldsymbol{O}_{p}(1)$ or $\boldsymbol{o}_{p}(1)$ if all it’s elements are ${O}_{p}(1)$ or ${o}_{p}(1)$ respectively.
@@ 2 Kernel Method
The kernel method uses an implicit mapping of the data ($X$) space into a higher-dimensional space [^1] This higher dimensional space is Hilbert space where the inner product of the vectors is well defined. ($\mathcal{F}$), which contains non-linear transformations of the given data. We denote this transformation by $\varphi:X\rightarrow\mathcal{F}$. This transformation offers a significant advantage since many non-linear relationships can be reformulated as linear relationships in an appropriately transformed space. Transformation to a higher-dimensional space using the kernel method makes a large set of non-linear forms available, rendering the discovery of a nonlinear relationship very likely. To illustrate how a nonlinear relation can be reformulated as a linear relation with appropriately transformed variables, consider the following example. We generate 2 variables $X$ and $Y$ from uniform distribution $U[-2,2]$. Define a binary variable $z$ as:
@@
It proves challenging if we attempt to use a linear boundary to separate the two classes of variable $z$. However, upon transforming the original spaces $X$ and $Y$ to $\varphi_{1}(X)=\sin^{2}(X)$ and $\varphi_{2}(Y)=\cos^{2}(Y)$ respectively, we find that the two classes can be easily distinguished as seen in figure-1.
@@
The blue points are in class 1, and the red ones are in class -1. Note that on the left-hand side, no line can achieve a decent accuracy in classification in the original space. However, we can easily do so in the transformed space on the right.
For the sake of simplicity, this example illustrates the transformation of a two-dimensional input $(X,Y)$ into a two-dimensional feature space. However, the transformed space $\varphi(X)$ is typically high-dimensional and potentially infinite-dimensional. As the dimensions of the transformed space grow, it captures the non-linearity better by virtue of the presence of more non-linear transformations of the input data. But this high (infinite)-dimensionality of transformed space seemingly makes computations extremely challenging (infeasible). The kernel method provides a solution to this issue. Our problem necessitates the computation of the similarity measure between transformed data points $\varphi(x_{i})$ and $\varphi(x_{j})$ via the inner product $\varphi(x_{i})^{\prime}\varphi(x_{j})$, which appears to require knowledge of the function $\varphi(\cdot)$. The kernel function $\mathcal{K}(\cdot,\cdot):X\times X\rightarrow R$ circumvents this requirement by computing the aforementioned similarity measure without necessitating knowledge of the feature map $\varphi(\cdot)$. Mercer’s condition ensures the existence of a feature map $\varphi(\cdot)$ for a valid (positive definite) kernel function in Hilbert space. A detailed discussion of Mercer’s condition can be found in Appendix - A.2. In the remaining part of this section, we illustrate how the kernel function computes the inner products by showing an explicit functional form of $\varphi(\cdot)$.
Some Popular Kernel Methods and Their Working
Polynomial Kernel
Then, the corresponding kernel function becomes:
$\mathcal{K}(a,b)=$ $\varphi(a)^{\prime}\varphi(b)$
Gaussian Kernel
This kernel goes beyond the high-dimensional kernel. This kernel is, in fact, infinite-dimensional. Let $\mathrm{x},\mathrm{z}\in\mathbb{R}^{k}$ and $\mathcal{K}(\mathrm{x},\mathrm{z})=e^{-\gamma\|\mathrm{x}-\mathrm{z}\|^{2}}$. Then, through the Taylor expansion, we can write
$=$ $\varphi(\mathrm{x})^{\prime}\varphi(\mathrm{z})$
@@ 3 The Estimator
Below, we list the data generation process for the transformed predictor set ($\varphi(\boldsymbol{x}_{t})$), the target ($\boldsymbol{y}$), and the proxies employed for supervision ($\boldsymbol{Z}$). Given the data structure, it is easy to explain why this supervised methodology is effective in estimating the target relevant factors.
@@
$\varphi$ maps our N-dimensional predictors $\boldsymbol{x}_{t}$ to a higher M-dimensional space. Assumption 1 endows this transformed set of predictors with a factor structure. An underlying factor structure among $\boldsymbol{X}$ implies the existence of a low dimensional plane, projection onto which explains maximal variation in the predictors. An equivalent interpretation of a linear factor structure on $\varphi(\boldsymbol{x}_{t})$ would be the existence of a lower dimensional manifold which explains maximum variation in $\boldsymbol{x}_{t}$. The basis of this manifold is comprised of a few uni-dimensional orthogonal projections of $\varphi(\boldsymbol{x}_{t})$.
At this juncture, it’s instructive to discuss the difference between our methodology and a set of non-linear dimension reduction methods categorized under the umbrella term ‘Sufficient Dimension Reduction (SDR)’. Sufficient dimension reduction techniques hinge upon the existence of a central subspace for $\boldsymbol{X}$, such that for every $t$, ${y}_{t+h}\perp\boldsymbol{x}_{t}|H^{T}\boldsymbol{x}_{t}$, where $H$ is an $N\times L$ matrix with $L<<N$, is a small fixed number. Matrix $H$ projects $\boldsymbol{x}_{t}$ onto an L-dimensional subspace, termed the central subspace. This projection is deemed ‘sufficient’ for predicting ${y}_{t+h}$, indicating that $\boldsymbol{x}_{t}-H^{T}\boldsymbol{x}_{t}$ is redundant from the perspective of predicting $y_{t+h}$. SDR accommodates non-linearity by permitting ${y}_{t+h}$ to be a non-linear function (called the link function) of $H^{T}\boldsymbol{x}_{t}$, i.e. SDR posits that
${y}_{t+h}=f(H^{T}\boldsymbol{x}_{t})\hskip 5.69054pt\text{for some unknown function}\hskip 2.84526ptf$
For more details, see Cook, (2018) and Li, (2018).
[…]
We now delineate the infeasible three-passes below.
@@ 3. Run time series regression of $y_{t+h}$ on predictive factors $\hat{\boldsymbol{F}}_{t}$,
Figure 2: Implementation of the Three Pass regression filter for the case T=4 and L=3 relevant factors. The variables $z^{(1)}\dots z^{(3)}$ and $f^{(1)}\dots f^{(3)}$ are the vectors representing the time series of the respective variables. $X_{s}$, (resp $\varphi(X_{s}))$ represents the cross section of $\boldsymbol{X}$ (resp $\varphi(\boldsymbol{X})$) in period $s$.
These three passes rely on the fact that the correlation between the transformed $\varphi(\boldsymbol{X})$ and the proxies is only due to target relevant factors. Therefore, pass 1 of the regression asymptotically yields a rotation of the relevant-factor loadings of the $j^{th}$ predictor. Cross sectional covariance between these loadings and the predictors, across $t$, is solely affected by the target relevant factor(s). Hence, pass 2 of this process traces the factor(s) out as a slope parameter. Although these three passes offer valuable insights into the mechanics of our process, they are infeasible in practise due to the unavailability of the transformed inputs $\varphi(\boldsymbol{X})$. This is where, the kernel trick shall prove to be handy. To see this, we note that factor(s), their predictive coefficients and the forecast can be expressed in closed form as below,
@@
$=\left(\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\boldsymbol{Z}\right)^{-1}\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\boldsymbol{Z}\left(\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\boldsymbol{Z}\right)^{-1}\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\boldsymbol{y}$
@@
These are obtained by simply replacing $\boldsymbol{X}$ by $\varphi(\boldsymbol{X})$ in the three-pass regression filter of Kelly & Pruitt, (2015). As evident from the expression of $\hat{\boldsymbol{F}}^{\prime}$, our three-pass filter applied on the transformed predictor space results in a favorable scenario where the eventual estimate of the factor(s) depends upon the transformed predictors only through their dot products in the transformed space. This holds true for $\hat{\boldsymbol{\beta}}$ and $\hat{\boldsymbol{y}}$ as well.
This inner product can be computed using a suitable kernel function. Alternatively, it can be inferred that employing a positive semidefinite (psd) kernel function to calculate dot products in these derived expressions is akin to executing the three-pass filter process on the implied transformed space, which, according to Mercer’s theorem, is guaranteed to exist.
The Kernel three-pass regression, just like the linear 3PRF, relies on the availability of suitable proxies. Kelly & Pruitt, (2015) shows that such proxies can always be constructed using the target variable $\boldsymbol{y}$. That process is explained below.
@@
    $\mathbb{E}\left\|\boldsymbol{F}_{t}\right\|^{4}<R,T^{-1}\sum_{s=1}^{T}\boldsymbol{F}_{s}\underset{T\rightarrow\infty}{\stackrel{{\scriptstyle p}}{{\longrightarrow}}}\boldsymbol{\mu}\text{ and }T^{-1/2}\left(\boldsymbol{F}^{\prime}\boldsymbol{J}_{T}\boldsymbol{F}-\boldsymbol{\Delta}_{F}\right)=\boldsymbol{O}_{p}(1)$.
@@
    $\mathbb{E}\left\|\phi_{i}\right\|^{4}\leq R$,$M^{-1}\sum_{j=1}^{N}\phi_{j}\underset{N\rightarrow\infty}{\stackrel{{\scriptstyle p}}{{\longrightarrow}}}\boldsymbol{0}$, $M^{-1/2}\left(\boldsymbol{\Phi}^{\prime}\boldsymbol{\Phi}-\mathcal{P}\right)=\boldsymbol{O}_{p}(1)$.
@@
Assumption 2.1 requires that our factors are regular in the sense that their covariance matrix is well-behaved asymptotically. Assumption 2.2 is an adaptation from Kelly & Pruitt, (2015). Since we assume a factor structure on the transformed space instead of the original predictor space, the normalization in various terms features $M$ and not $N$, where $M$ is the dimension of our transformed space. Assumptions 2.3-2.5 borrowed from Kelly & Pruitt, (2015), impose regularity on various error processes.

### v3

@@ Abstract
We forecast a single time series using a high-dimensional set of predictors. When predictors share common underlying dynamics, a latent factor model estimated by the Principal Component method effectively characterizes their co-movements. These latent factors succinctly summarize the data and aid in prediction, mitigating the curse of dimensionality. However, two significant drawbacks arise: (1) not all factors may be relevant, and utilizing all of them in constructing forecasts leads to inefficiency, and (2) typical models assume a linear dependence of the target on the set of predictors, which limits accuracy. We address these issues through a novel method: Kernel Three-Pass Regression Filter. This method extends a supervised forecasting technique, the Three-Pass Regression Filter, to exclude irrelevant information and operate within an enhanced framework capable of handling nonlinear dependencies. Our computationally efficient method demonstrates strong empirical performance, particularly over longer forecast horizons.
Keywords: Forecasting, High dimension, Approximate factor model, Reproducing Kernel Hilbert space, Three-pass regression filter, Machine Learning.
@@ 1 Introduction
In recent years, the surge in high-dimensional datasets across fields like economics has ushered in new opportunities and challenges. A paramount issue is the ‘curse of dimensionality,’ which undermines the effectiveness of traditional finite-dimensional estimation methods. Most modeling techniques applied to high-dimensional data assume the existence of a low-dimensional structure that effectively summarizes the data. One stylized feature of high-dimensional economic datasets is the presence of high and pervasive collinearity among variables, leading researchers to posit a data-generating process that assumes all variables are a function of a few latent factors. This formulation is commonly referred to as the factor model. A vast amount of literature focuses on using this latent factor structure for forecasting applications. A typical example is found in diffusion index models (Stock & Watson, (2002b)), where latent factors are derived from a high-dimensional set of variables using Principal Components (hereafter, PC) method. These factors are subsequently utilized to forecast a target variable. A limitation of this PC-based factor estimation is its unsupervised nature, i.e., no information from the target variable is incorporated.
@@
Since the primary goal is to forecast a target rather than estimate the underlying factor structure, introducing a degree of supervision can be beneficial. This can help filter out irrelevant information from the predictor set, thus enhancing the predictive accuracy. This can be done in different ways: using soft and hard thresholding methods to remove predictors with no predictive content, as in Bai & Ng, (2008), or assigning varying weights to predictors based on their predictive capabilities for the target (see, for example, Huang et al., (2022)), or estimate the subset of factors that exhibit predictive power for the target rather than the complete set of factors that drive the predictors, as in Kelly & Pruitt, (2015).
@@
The aforementioned models, whether utilizing PCA or supervised factor models, are predicated on the convenient assumption of linearity. However, as underscored in Goulet Coulombe et al., (2022), non-linearity often characterizes many predictive relationships in economics, particularly over extended time horizons and within data-rich environments.
@@
Various approaches have been proposed to integrate non-linearity into factor models. For instance, squared principal components (PCs) or principal component squared ($PC^{2}$) as seen in Bai & Ng, (2008), sufficient forecasting by Fan et al., (2017), the kernel trick to estimate factors (Kutateladze, (2022)) among others. However, these approaches have limited supervision in the prediction process, if any. For example, Fan et al., (2017) estimates factors through an unsupervised method (PC) and then derives sufficient indices using these PCs. Similarly, Kutateladze, (2022) essentially applies kernel PCA (an unsupervised method) to estimate the set of factors driving a higher-dimensional space obtained by lifting the set of predictors through the kernel method. In Bai & Ng, (2008), a very particular form of non-linearity (quadratic) is examined, which is somewhat ad hoc. Although they employ thresholding methods to reduce predictors to a smaller set, their screening method, however, may still encounter challenges in filtering relevant factors within this subset, leading to inefficient forecasts.
@@
Our paper incorporates both non-linearity and supervision by introducing a novel kernel three-pass regression filter. Our approach essentially applies the three-pass filter (hereafter 3PRF) proposed by Kelly & Pruitt, (2015) to a transformed set of predictors. We adopt the lifting concept similar to Kutateladze, (2022), but instead of employing an unsupervised method like kernel PCA, we utilize a supervised method to estimate factors relevant to the target variable.
@@
The table below summarizes our discussion by listing some popular methods [^1] The entries in this table are some of the most popular forecasting methods used in econometric literature. However, by no means do they form an exhaustive set. in literature and how this paper is placed among them
@@
Table 1: Factor Model Based Forecasting Methods
The paper proceeds as follows. Section 2 provides a brief introduction to Kernel methods. Section 3 introduces our estimator and discusses its similarity with the estimator of Kelly & Pruitt, (2015). We also list a set of assumptions that ensure the theoretical properties of our estimators, which are given in the subsequent section 4. We present our empirical results in sections 5 and 6 and conclude in section 7. Mathematical proofs and implementation details are given in the appendix.
@@
We use $\boldsymbol{y}$ to denote the $\text{T}\times 1$ vector of the target variable, i.e. $\boldsymbol{y}=(y_{h},y_{h+1}\dots y_{t+h})$. We have $N$ predictors with $T$ observations for each predictor. The cross section of predictors at a time $t$ is given by the $N\times 1$ vector $\boldsymbol{x}_{t}$. Similarly, the vector of temporal observations of a predictor $i$ is given by $\mathbf{x}_{i}$. We stack the predictors in a ${T}\times{N}$ matrix $\boldsymbol{X}$, $\boldsymbol{X}=$ $\left(\boldsymbol{x}_{1}^{\prime},\boldsymbol{x}_{2}^{\prime},\ldots,\boldsymbol{x}_{T}^{\prime}\right)^{\prime}=\left(\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{N}\right)$. We have $L$ proxies which we stack in a ${T}\times\ {L}$ matrix $\boldsymbol{Z}=\left(\boldsymbol{z}_{1}^{\prime},\boldsymbol{z}_{2}^{\prime},\ldots,\boldsymbol{z}_{T}^{\prime}\right)^{\prime}$. The demeaning matrix $\boldsymbol{J}_{T}\equiv\boldsymbol{I}_{T}-\frac{1}{T}\mathbf{\iota}_{T}\mathbf{\iota}_{T}^{\prime}$, where $\boldsymbol{I}_{T}$ is the $T$-dimensional identity matrix and $\mathbf{\iota}_{T}$ the $T$-vector of ones. For matrices $\boldsymbol{U}$ and $\boldsymbol{V}$ of conformable dimensions, $\boldsymbol{S}_{UV}\equiv\boldsymbol{U}^{\prime}\boldsymbol{J}_{T}\boldsymbol{V}$. For the transformed set of predictors $\varphi(\boldsymbol{X})$, $\varphi_{j}(\mathbf{x})$ denotes the $j^{th}$ observation. $\varphi(\boldsymbol{X})=$ $\left(\varphi_{1}(\mathbf{x}),\varphi_{2}(\mathbf{x}),\ldots,\varphi_{M}(\mathbf{x})\right)=$ $\left(\varphi(\boldsymbol{x}_{1})^{\prime},\varphi(\boldsymbol{x}_{2})^{\prime},\ldots,\varphi(\boldsymbol{x}_{T})^{\prime}\right)^{\prime}$. Stochastic orders are denoted by the usual ${O}_{p}$ and ${o}_{p}$. For a matrix, $\boldsymbol{O}_{p}$ and $\boldsymbol{o}_{p}$ denotes the element wise stochastic order, i.e., a matrix is said to be $\boldsymbol{O}_{p}(1)$ or $\boldsymbol{o}_{p}(1)$ if all it’s elements are ${O}_{p}(1)$ or ${o}_{p}(1)$ respectively.
@@ 2 Kernel Method
Let $\varphi:\boldsymbol{X}\rightarrow\mathcal{F}$ denote a transformation of the original data into a higher-dimensional space [^2] Precisely, we are referring to a Hilbert space where the inner product of the vectors is well-defined. containing the original set of predictors and their non-linear transformations. Methods such as principal components or the three-pass regression filter depend on the input $\boldsymbol{X}$ only through the $T\times T$ matrix of dot products $\boldsymbol{X}^{\prime}\boldsymbol{X}$. Applying these methods to the transformed predictors $\varphi(\boldsymbol{X})$ would therefore require computing the inner product $\varphi(\boldsymbol{X})^{\prime}\varphi(\boldsymbol{X})$. This computation can be cumbersome or infeasible [^3] When the transformed space is infinite-dimensional. Here, the kernel trick proves to be handy, allowing us to calculate inner products within the transformed space without requiring explicit knowledge of $\varphi$. A valid [^4] A positive semi-definite kernel as discussed in Appendix-A.2. kernel corresponds to an inner product of features $\varphi(\boldsymbol{X})$, where the analytical form of the function $\varphi(\cdot)$ may be unknown, but it is guaranteed to exist by Mercer’s theorem (Appendix-A.2). Hence, using a Kernel function to compute inner products within a method is akin to performing the estimation exercise(implicitly) on the set of transformed features. In the Supplementary appendix-B.1, we illustrate how different kernel functions correspond to the inner products of transformed inputs.
Utilizing a transformed set of predictors provides a significant advantage, as many nonlinear relationships can be reformulated as linear relationships in the appropriately transformed space. As an illustration, consider the following example. We generate two variables $X$ and $Y$ from uniform distribution $U[-2,2]$. Define a binary variable $z$ as:
@@
As shown in the figure-1 (left), a linear boundary cannot separate the two classes of variable $z$. However, upon transforming the original spaces $X$ and $Y$ to $\varphi_{1}(X)=\sin^{2}(X)$ and $\varphi_{2}(Y)=\cos^{2}(Y)$ respectively, we find that the two classes can be easily distinguished as seen in figure-1(right). The blue points are in class 1, and the red ones are in class -1.
@@
For the sake of simplicity, this example illustrates the transformation of a two-dimensional input $\boldsymbol{W}=(X,Y)$ into a two-dimensional feature space. The transformed space $\varphi(\boldsymbol{W})$ is typically high-dimensional and potentially infinite-dimensional. Transformation to a higher-dimensional space makes a large set of non-linear forms available, rendering the discovery of a nonlinear relationship very likely.
@@ 3 The Estimator
Below, we list the data generation process for the transformed predictor set ($\varphi(\boldsymbol{X})$), the target ($\boldsymbol{y}$), and the proxies employed for supervision ($\boldsymbol{Z}$). Given the data structure, it is easy to explain why this supervised methodology is effective in estimating the target relevant factors.
@@
$\varphi$ maps our N-dimensional predictors $\boldsymbol{x}_{t}$ to a higher M-dimensional space. Assumption 1 endows this transformed set of predictors with a factor structure. An underlying factor structure among $\boldsymbol{X}$ implies the existence of a low dimensional plane, projection onto which explains maximal variation in the predictors. An equivalent interpretation of a linear factor structure on $\varphi(\boldsymbol{x}_{t})$ would be the existence of a lower dimensional manifold which explains maximum variation in $\boldsymbol{x}_{t}$. This manifold’s basis comprises a few uni-dimensional orthogonal projections of $\varphi(\boldsymbol{x}_{t})$.
The infeasible three-passes are summarized in 2 below.
@@ 3. Run time series regression of $y_{t+h}$ on predictive factors $\hat{\boldsymbol{F}}_{t}$,
Figure 2: Implementation of the Kernel Three Pass regression filter for the case T=4 and L=3 relevant factors. The variables $z^{(1)}\dots z^{(3)}$ and $f^{(1)}\dots f^{(3)}$ are the vectors representing the time series of the respective variables. $X_{s}$, (resp $\varphi(X_{s}))$ represents the cross section of $\boldsymbol{X}$ (resp $\varphi(\boldsymbol{X})$) in period $s$.
These three passes rely on the fact that the correlation between the transformed $\varphi(\boldsymbol{X})$ and the proxies is only due to target relevant factors. Therefore, pass 1 of the regression asymptotically yields a rotation of the relevant-factor loadings of the $j^{th}$ predictor. Cross-sectional covariance between these loadings and the predictors, across $t$, is solely affected by the target relevant factor(s). Hence, pass 2 of this process traces the factor(s) out as a slope parameter. The last pass involves regressing the target variable on the estimated factor(s). Although these three passes offer valuable insights into the mechanics of our process, they are infeasible in practice due to the unavailability of the transformed inputs $\varphi(\boldsymbol{X})$. This is where the kernel trick proves to be useful. To see this, we note that factor(s), their predictive coefficients, and the forecast can be expressed in closed form as below,
@@
$=\left(\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\boldsymbol{Z}\right)^{-1}\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\boldsymbol{Z}\left(\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\boldsymbol{Z}\right)^{-1}\times$
$\hskip 28.45274pt\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\boldsymbol{y}$
@@
These expressions are obtained by simply replacing $\boldsymbol{X}$ by $\varphi(\boldsymbol{X})$ in the three-pass regression filter of Kelly & Pruitt, (2015). As evident from the expression of $\hat{\boldsymbol{F}}^{\prime}$, the filtration process applied on the transformed predictor space results in a favorable scenario where the eventual estimate of the factor(s) depends upon the transformed predictors only through their dot products in the transformed space. This holds true for $\hat{\boldsymbol{\beta}}$ and $\hat{\boldsymbol{y}}$ as well.
This inner product can be computed using a suitable kernel function. Alternatively, it can be inferred that employing a positive semidefinite (psd) kernel function to calculate dot products in these derived expressions is akin to executing the three-pass filter process on the transformed set of predictor(s), which, according to Mercer’s theorem, are guaranteed to exist.
The Kernel three-pass regression, like the linear 3PRF, relies on the availability of suitable proxies. Kelly & Pruitt, (2015) show that such proxies can always be constructed using the target variable $\boldsymbol{y}$. That process is explained in table-3 below.
@@
    $\mathbb{E}\left\|\boldsymbol{F}_{t}\right\|^{4}<R,T^{-1}\sum_{s=1}^{T}\boldsymbol{F}_{s}\underset{T\rightarrow\infty}{\stackrel{{\scriptstyle p}}{{\longrightarrow}}}\boldsymbol{\mu}\text{ and }T^{1/2}\left(\dfrac{\boldsymbol{F}^{\prime}\boldsymbol{J}_{T}\boldsymbol{F}}{T}-\boldsymbol{\Delta}_{F}\right)=\boldsymbol{O}_{p}(1)$.
@@
    $\mathbb{E}\left\|\phi_{i}\right\|^{4}\leq R$,$M^{-1}\sum_{j=1}^{M}\phi_{j}\underset{N\rightarrow\infty}{\stackrel{{\scriptstyle p}}{{\longrightarrow}}}\boldsymbol{0}$, $M^{1/2}\left(\dfrac{\boldsymbol{\Phi}^{\prime}\boldsymbol{\Phi}}{M}-\mathcal{P}\right)=\boldsymbol{O}_{p}(1)$.
@@
Assumption 2.1 requires that our factors are regular in the sense that their covariance matrix is well-behaved asymptotically. Assumption 2.2 is an adaptation from Kelly & Pruitt, (2015). Since we assume a factor structure on the transformed space instead of the original predictor space, the normalization in various terms features $M$ and not $N$, where $M$ is the dimension of our transformed space. Assumptions 2.3-2.5, borrowed from Kelly & Pruitt, (2015), impose regularity on various error processes.